OpenStudy (anonymous):

3/2x-1/2x=0 make a graph?

OpenStudy (anonymous):

Well (3/2-1/2)x=0 x=0 so it is just the y-axis

OpenStudy (anonymous):

sory you simplification is wrong,, i have done it in the same way but this is may not be the answer

OpenStudy (anonymous):

Is it (3/2)x? Or 3/(2x) HUGE difference.

OpenStudy (anonymous):

Is it this? \[\frac{3}{2x} - \frac{1}{2x} = 0\]

OpenStudy (anonymous):

Xx0=1 see this simplification

OpenStudy (anonymous):

3/2x

OpenStudy (anonymous):

\[\frac{3}{2x} \] or \[\frac{3x}{2}\] ?

OpenStudy (anonymous):

3/2x-1/2x=0

OpenStudy (anonymous):

Its a hyperbola and it NEVER equals zero. So your problem is messed up.

OpenStudy (anonymous):

cyde, do you not understand the difference of (3x)/2 and 3/(2x)?

OpenStudy (anonymous):

cyde, click on the equation button and type in frac{}{} and put the numerator in the first set of curly braces and the denominator in the second set of curly braces

OpenStudy (anonymous):

3/(2x)-1/(2x)=0 graph the equation and find solution set?

OpenStudy (anonymous):

Its a hyperbola. 1/x((3/2)-(1/2))=y y=1/x It is a standard ellipse, it is true for all values of y and x but zero. \[\left\{ x \in \mathbb{R}|x \neq 0 \right\} \] For domain and \[\left\{ y \in \mathbb{R}|y \neq 0\right\}\]

OpenStudy (anonymous):

For the range^^